OpenStudy (anonymous):

The Ward family is redecorating several rooms of their house. They decided to use carpet tiles in the family room. The room has an area of 176 square feet and is 5 feet longer than it is wide. Find the dimensions of the family room.

OpenStudy (anonymous):

L=w+5 w(w+5)=176 w^2+5w=176 w^2+5w-176 w=11 L=w+5 L=11+5 16 Length=16 Width=11

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (anonymous):

you're welcome. :)